How do you solve for substitution equations?
-4x + y = 6
-5x - y = 21

Found 2 solutions by stanbon, MathTherapy:
Answer by stanbon(75887)

How do you solve for substitution equations?
-4x + y = 6
-5x - y = 21
----
Solve one of the equations for "y"::
y = 4x+6
---
Substitute for "y" in the other equation and solve for "x"::
-5x -4x-6 = 21
------
-9x = 27
-----
x = -3
------------
Solve for "y"::
y = 4x+6
y = 4*-3+6
y = -6
-------------
cheers,
Stan H.
-------------

Answer by MathTherapy(10858)


How do you solve for substitution equations?
-4x + y = 6
-5x - y = 21
Do you really want to solve by substitution? The equations can be solved by ELIMINATING the y variable instantly, 
without having to change any of the 2.